# TokenWeave: Efficient Compute-Communication Overlap for Distributed LLM Inference

### v2

@@
We present TokenWeave to address these challenges. TokenWeave proposes a Token-Splitting technique that divides the tokens in the inference batch into two approximately equal subsets in a wave-aware manner. The communication of one subset is then overlapped with the computation of the other. In addition, TokenWeave optimizes the order of the layer normalization computation with respect to communication operations and implements a novel fused AllReduce–RMSNorm111In all modern implementations, residual addition is fused with RMSNorm, and all our results use this fused implementation. For brevity, we mention only RMSNorm in all our discussions. kernel that carefully leverages Multimem instruction support available on NVIDIA Hopper GPUs. These optimizations allow TokenWeave to perform communication and RMSNorm using only SMs. Moreover, our kernel enables the memory-bound RMSNorm to be overlapped with the other batch’s computation, providing additional gains.
@@ 2.2. Compute-Communication Overlap
Performance comparison of AllReduce vs ReduceScatter plus AllGather on H100 GPUs for a hidden size of .
Performance comparison of AllReduce vs ReduceScatter plus AllGather on H100 GPUs for a hidden size of .
@@ 2.3. NV-SHARP, Multicast and SymmetricMemory
PyTorch exposes NVLS through its SymmetricMemory API (Wang et al., 2025). SymmetricMemory facilitates allocating peer buffers on GPUs using the symm_mem.empty
call (similar to torch.empty()
). After buffer allocation, a collective call to symm_mem.rendezvous()
exchanges memory handles, mapping peer buffers into the virtual address space of each participating GPU. Post-rendezvous, each GPU can access remote or multicast pointers using standard memory operations within Triton or CUDA kernels, eliminating explicit NCCL calls and substantially simplifying the implementation of communication routines.
@@ 3. TokenWeave: Design and Implementation
Two diagrams: (a) shows compute and communication without overlap; (b) shows them overlapping.
@@
dualAdd–RMSNorm reordering, and a fused AllReduce–RMSNorm implementation. Figure 8(b) provides a high-level schematic of our method compared to the standard tensor-parallel implementation.
@@ 3.1.2. Overlapped Execution
The operations of these two batches can now be overlapped as shown in Figure 8(b). As shown, when the AllReduce of the first batch is being processed, we compute the attention of the second batch. The FFN of the first batch is then overlapped with the AllReduce of the second batch and so on. We implement this overlapped execution via CUDA streams. A communication stream handles all the communication operations, while a compute stream runs the compute operations. Lightweight synchronization using CudaStreamSynchronize is performed between the compute and communication stream to handle data dependencies.
@@ 4.1. Experimental Setup
Environment: All our experiments are performed on an H100 NVIDIA DGX system with NV-SHARP support. The experiments use the same system but use four of the eight GPUs. We also use fixed TDP settings to ensure stable performance measurements (NVIDIA, 2020). For microbenchmarks, we perform warm-up, flush the L2 cache and use cudaEvent
s for timing measurements. For end-to-end (e2e) results, we perform warm-up and then use time.perf_counter()
for timing.
@@
We also compare with TileLink (Zheng et al., 2025) which is the current state-of-the-art method for reducing communication overheads using compute-communication fusion and outperforms Flux (Chang et al., 2024), as well as NanoFlow (Zhu et al., 2024), a technique that optimizes compute, memory-bandwidth and communication. However, note that TileLink is not integrated with vLLM which makes an end-to-end comparison difficult. We thus use optimized single layer implementations of various models to compare TokenWeave against TileLink. NanoFlow again has their own serving stack and is not integrated with vLLM. Thus, to compare against the NanoFlow baseline, we look at the communication overhead which is recovered in NanoFlow against their baseline framework vs that recovered by TokenWeave against vLLM, which is our baseline framework. We also report absolute numbers for reference.
@@ 4.2.1. TokenWeave Latency Benefits
Two diagrams: (a) shows compute and communication without overlap; (b) shows them overlapping.
@@ 4.2.2. TokenWeave Throughput Gains
Figure 12(b) shows the performance evaluation with chunked-prefills on workloads with both fixed input and output length requests, as well as ShareGPT (Wang et al., 2023) which is a real-world workload trace with varying input/output requests lengths based on user conversations. As shown, in the H100 case, TokenWeave continues to show substantial throughput gains of approximately for dense models, effectively recovering most of the communication overhead.
@@ 4.2.4. Comparison with NanoFlow
We evaluate the adapted NanoFlow implementation under two conditions: with NanoFlow enabled and disabled within its own serving stack. Figure 15 shows the performance evaluation with chunked-prefills (chunk size ) on workloads with various fixed input and output length requests. As shown, NanoFlow provides only modest performance gains in the range of of to , closely matching the communication improvement reported in the NanoFlow paper. In contrast, TokenWeave demonstrates significantly larger improvements of approximately across a broad range of scenarios (see Figure 12(b)).

### v4

@@
We present TokenWeave to address these challenges. TokenWeave proposes a Token-Splitting technique that divides the tokens in the inference batch into two approximately equal subsets in a wave-aware manner. The communication of one subset is then overlapped with the computation of the other. In addition, TokenWeave optimizes the order of the layer normalization computation with respect to communication operations and implements a novel fused AllReduce–RMSNorm111In all modern implementations, residual addition is fused with RMSNorm, and all our results use this fused implementation. For brevity, we mention only RMSNorm in all our discussions. kernel that carefully leverages Multimem instruction support available on Hopper and Blackwell NVIDIA GPUs. These optimizations allow TokenWeave to perform communication and RMSNorm using only SMs. Moreover, our kernel enables the memory-bound RMSNorm to be overlapped with the other batch’s computation, providing additional gains.
@@ 2.3. NV-SHARP, Multicast and SymmetricMemory
PyTorch exposes NVLS through its SymmetricMemory API (Wang et al., 2025). SymmetricMemory facilitates allocating peer buffers on GPUs using the symm_mem.empty call (similar to torch.empty()). After buffer allocation, a collective call to symm_mem.rendezvous() exchanges memory handles, mapping peer buffers into the virtual address space of each participating GPU. Post-rendezvous, each GPU can access remote or multicast pointers using standard memory operations within Triton or CUDA kernels, eliminating explicit NCCL calls and substantially simplifying the implementation of communication routines.
@@ 3. TokenWeave: Design and Implementation
dualAdd–RMSNorm reordering, and a fused AllReduce–RMSNorm implementation. Figure 8 provides a high-level schematic of our method compared to the standard tensor-parallel implementation.
@@ 3.1.2. Overlapped Execution
The operations of these two batches can now be overlapped as shown in Figure 8. As shown, when the AllReduce of the first batch is being processed, we compute the attention of the second batch. The FFN of the first batch is then overlapped with the AllReduce of the second batch and so on. We implement this overlapped execution via CUDA streams. A communication stream handles all the communication operations, while a compute stream runs the compute operations. Lightweight synchronization using torch.cuda.stream_wait(stream, wait_stream) (it doesn’t involve cpu) is performed between the compute and communication stream to handle data dependencies.
@@ 4.1. Experimental Setup
Environment: All our experiments are performed on an H100 NVIDIA DGX system with NV-SHARP support. The experiments use the same system but use four of the eight GPUs. We also use fixed TDP settings to ensure stable performance measurements (NVIDIA, 2020). For microbenchmarks, we perform warm-up, flush the L2 cache and use cudaEvents for timing measurements. For end-to-end (e2e) results, we perform warm-up and then use time.perf_counter() for timing.
@@
We also compare with TileLink (Zheng et al., 2025) which is the current state-of-the-art method for reducing communication overheads using compute-communication fusion and outperforms Flux (Chang et al., 2024), as well as NanoFlow (Zhu et al., 2024), a technique that optimizes compute, memory-bandwidth and communication. However, note that TileLink222https://github.com/ByteDance-Seed/Triton-distributed/tree/b10fa2e is not integrated with vLLM which makes an end-to-end comparison difficult. We thus use optimized single layer implementations of various models to compare TokenWeave against TileLink. NanoFlow333https://github.com/efeslab/Nanoflow/tree/e2e3c2a again has their own serving stack and is not integrated with vLLM. Thus, to compare against the NanoFlow baseline, we look at the communication overhead which is recovered in NanoFlow against their baseline framework vs that recovered by TokenWeave against vLLM, which is our baseline framework. We also report absolute numbers for reference.
@@ 4.2.2. TokenWeave Throughput Gains
Figure 12 shows the performance evaluation with chunked-prefills on workloads with both fixed input and output length requests, as well as ShareGPT (Wang et al., 2023) which is a real-world workload trace with varying input/output requests lengths based on user conversations. As shown, in the H100 case, TokenWeave continues to show substantial throughput gains of approximately for dense models, effectively recovering most of the communication overhead.
@@ 4.2.4. Comparison with NanoFlow
We evaluate the adapted NanoFlow implementation under two conditions: with NanoFlow enabled and disabled within its own serving stack. Figure 15 shows the performance evaluation with chunked-prefills (chunk size ) on workloads with various fixed input and output length requests. As shown, NanoFlow provides only modest performance gains in the range of of to , closely matching the communication improvement reported in the NanoFlow paper. In contrast, TokenWeave demonstrates significantly larger improvements of approximately across a broad range of scenarios (see Figure 12).